

********************************************************************
UARS WEEKLY STATUS REPORT - 18 AUG 2000
********************************************************************
***********
SOLSTICE
***********
SOLSTICE is in good health, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Our solar data is not affected by the larger than normal
pointing errors since they are within our calibrations. Additionally, we
continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
Level2 processing is halted due to bugs in the OA_SSPP_ATT service. CDHF
and CSC personnel are working on the problem.
***********
SUSIM
***********
Except for the turnoff for the recent yaw-around and the subsequent
warmup period, SUSIM continues to gather a normal set of daily solar
scans scheduled for periods of predicted telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in a meeting held to discuss the status
of calibration efforts for the UARS Fixed Head Star Tracker 1.
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinate with the Flight
Operations Team and to assist the Project Scientist in the
scheduling of instrument operations.
***********
CDHF
***********
New UOAS engineering delivery made at SOLSTICE request is available.
(oa_sspp_att services)
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
5 days SEN (5.0) PROD (1992) (5R-5C=0L)
1 day SUS (5.0) TEST (1R-1C=0L)
Data created in the month of June 2000 has been previewed for HALOE, NMC,
and UKMO in preparation for the DAAC.
UARS CDHF REHOST:
- UARS PCG has notified all PI Representatives regarding plan to run in
a parallel mode on the current and rehost systems for at least (2) weeks
starting 31-Jul-2000. We are planning to extend test another week.
Investigating SUSIM run which did not compare correctly.
***********
FOT
***********
REPORTING PERIOD
This report covers 05 August 2000 (Orbit 48647, GMT Day 218) through 11 August
2000 (Orbit 48751, GMT Day 224).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
AXIS-1 218/2357 48661 Off
WINDII 222/1244 48714 Off
HALOE 223/2304 48736 Off
The beta angle ranged between the angles of 48.3 degrees to 13.4 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
+2 degrees on 14 August 2000 (DOY 227).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.5 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 04 August 2000 and 11 August 2000 are provided in Appendix D.
The clock error ranged from -12.9 to 14.3 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 5
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 37
¥ WINDII = 01
¥ HALOE = 01
¥ PEM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 05
Five (5) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/153 - 2287.496961
· Transponder B: 00/213 - 2287.499288
GROUND SYSTEM OPERATION
Five (5) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/223 show the observatory data loss to be 297 hours, 50
minutes, 22.560 seconds (an increase of 50 minutes 22.560 seconds since the
last report period). This is a 0.40 percent data loss which equals a 99.60
percent data capture for the mission.
The increase of 50 minutes 22.560 seconds occurred over one (1) day:
· 50 minutes 22.560 seconds on Day 221 due to no acquisition.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to operate FHST-1 for approximately one orbit each day. By
cycling the tracker on and off for such a short duration the degradation seems
to have been reduced significantly. Scale factor calibrations are being
provided on a regular basis from FDF and are being uplinked when received. It
appears that the FOT has been able to stabilize the performance of FHST-1 for
now and continues to work on establishing another viable attitude control
solution should the tracker fail.
UARS had an anomaly this week which resulted in a 50 minute data loss. This
data loss was a result of the wrong time tags being sent for a segment of the
daily ATC load. This mistake caused the stored command processor to halt on
the first location of this load and thus all of the subsequent commands were
not executed. Among the commands not executed were the HGA slew commands which
slew the antenna from one TDRS to the other. The loss of these commands
stopped the HGA and thus data was lost until the problem was isolated and
corrected.
MEETINGS
The FOT attended a meeting to discuss FHST-1 on 10 August 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for August 14, 2000 (DOY 227). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
_______________________________________________________________________________
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 11 Aug 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
48650 218/0537 00-389 NBTR NOOPS D-218
48656 218/1558 00-389 NBTR NOOPS D-218
48656 218/1558 00-391 FHST-1 POWER ON
48657 218/1738 00-391 FHST-1 POWER OFF
48658 218/1840 00-392 NBTR NOOPS D-219
48661 218/2357 00-393 AXIS-1 POWER OFF
48664 219/0457 00-392 NBTR NOOPS D-219
48669 219/1237 00-392 NBTR NOOPS D-219
48671 219/1600 00-391 FHST-1 POWER ON
48672 219/1739 00-391 FHST-1 POWER OFF
48673 219/1844 00-394 NBTR NOOPS D-220
48681 220/0725 00-395 OBC Clk Adj 0.0 to 20.6 msec/day
48685 220/1421 00-394 NBTR NOOPS D-220
48686 220/1601 00-397 FHST-1 POWER ON
48687 220/1741 00-397 FHST-1 POWER OFF
48688 220/1922 00-396 NBTR NOOPS D-221
48689 220/2035 00-398 FHST-1 SCALE FACTOR UPDATE
48696 221/0649 00-396 NBTR NOOPS D-221
48697 221/0917 00-399 OBC Clk Adj 20.6 to 0.00 msec/day
48701 221/1520 00-400 FHST-1 ALIGNMENT CALIBRATION
48701 221/1603 00-400 FHST-1 POWER ON
48702 221/1742 00-400 FHST-1 POWER OFF
48706 222/0004 00-401 NBTR NOOPS D-222
48714 222/1244 00-403 WINDII POWER OFF
48716 222/1522 00-401 NBTR NOOPS D-222
48716 222/1604 00-404 FHST-1 POWER ON
48717 222/1744 00-404 FHST-1 POWER OFF
48717 222/1744 00-406 FHST-1 SCALE FACTOR UPDATE
48718 222/1927 00-405 NBTR NOOPS D-223
48726 223/0656 00-405 NBTR NOOPS D-223
48729 223/1246 00-407 OBC Clk Adj 0.0 to -20.6 msec/day
48731 223/1606 00-404 FHST-1 POWER ON
48731 223/1606 00-405 NBTR NOOPS D-223
48732 223/1749 00-404 FHST-1 POWER OFF
48736 223/2304 00-408 HALOE POWER OFF
48743 224/1025 00-409 NBTR PLAYBACK
48746 224/1527 00-410 FHST-1 SCALE FACTOR UPDATE
48746 224/1607 00-404 FHST-1 POWER ON
48748 224/1901 00-411 OBC Clk Adj -20.6 to 0.0 msec/day
48748 224/1901 00-404 FHST-1 POWER OFF
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 11 Aug 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
OBSVLBEON Y-LO LOAD BUS VOLTS EON 48714
PWLBV Y-LO LOAD BUS VOLTAGE 48714
OBSBATVEON2 Y-LO BAT 2 VOLT EON 48714
OBSBATVEON3 Y-LO BAT 3 VOLT EON 48714
PWBAT3V Y-LO BAT 3 VOLTAGE 48714
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 11 Aug 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
48673 219/07:28 (06 Aug 00) from 0.00 to 20.6 msec/day
48697 221/09:20 (08 Aug 00) from 20.6 to 0.00 msec/day
48729 223/12:50 (10 Aug 00) from 0.00 to -20.6 msec/day
48748 224/19:10 (11 Aug 00) from -20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 11 Aug 2000
04 Aug 2000 - GMT Day 217 Beta = 48.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- --------- ---------- ------ ------- ------
1 +78.4/+56.0 -3.05/-4.31 +0.0/+0.0 15.5 0.000 0.0
2 +44.8/-33.6 +5.79/+4.19 +27.2/-15.2 25.0 1.020 21.6
3 +22.4/-44.8 +4.51/+3.23 +26.8/-15.2 24.8 1.020 21.2
11 Aug 2000 - GMT Day 224 Beta = 13.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------ ------- ------
1 +89.6/+67.2 -3.05/-4.62 +0.0/+0.0 15.5 0.000 0.0
2 +16.8/-22.4 +5.47/+3.87 +29.2/-13.2 26.2 1.020 20.3
3 +22.4/-22.4 +4.51/+2.60 +29.4/-13.2 26.2 1.018 20.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 11 Aug 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-035 FOT 48649 218/0455 22770 Drop outs/Ret. Link
00-036 FOT 48672 219/1739 22775 Drop outs
00-037 FOT 48685 220/1338 14170 Generic Late Acq.
00-038 FOT 48701 221/1520 N/A Prime Pack Crashed
00-039 FOT 48703 221/1924 22785 Wrong ATC Loaded
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-026 FOT see list June 00 14170 Generic Late Acq.
00-031 FOT see list July 00 14170 Generic Late Acq.
00-034 FOT 48566 212/1547 22739 No Lock/Bad Vector
00-035 FOT 48649 218/0455 22770 Drop outs/Ret. Link
00-036 FOT 48672 219/1739 22775 Drop outs
00-038 FOT 48701 221/1520 N/A Prime Pack Crashed
00-039 FOT 48703 221/1924 22785 Wrong ATC Loaded
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-037 FOT see list Aug 00 14170 Generic Late Acq.
00-032 FHST 48211 188/2242 N/A FHST 2 Anomaly
SEVNTFW00/32